# Model Predictive Control via Probabilistic Inference: A Tutorial and Survey

### v2

@@ Abstract
This paper provides a tutorial and a survey of probabilistic inference-based model predictive control (PI-MPC) for robotics. PI-MPC defines an optimal control distribution shaped by trajectory cost, a control prior, and a temperature parameter. In the tutorial part, we derive this view and describe action generation via variational inference, highlighting Model Predictive Path Integral (MPPI) control as a representative algorithm. In the survey part, we organize the PI-MPC literature around the principal design choices identified in the tutorial: prior distribution design, multi-modal distribution handling, constraint satisfaction, scalability to high-degree-of-freedom robots, hardware acceleration, and theoretical foundations. Overall, this paper aims to serve as a coherent entry point for researchers and practitioners interested in understanding, implementing, and extending PI-MPC.
@@
articletype: REVIEW ARTICLE
@@ 1 Introduction
Establishing a universal control theory applicable to diverse robots and tasks has long been a fundamental goal in robotics research. Model Predictive Control (MPC) offers an intuitive and general framework for robot control by optimizing actions while predicting future states over a finite horizon. To apply MPC to real-world robots and tasks, we may need to formulate complex optimal control problems that capture the intricacies of robot dynamics and task requirements. However, these formulations often exhibit strong nonlinearity, making them intractable for traditional numerical optimization methods widely familiar in the control field. In many robotic applications, what we need is not necessarily strong theoretical guarantees (e.g., infinite-horizon stability), but an easy-to-use optimization framework that can reliably solve a broad class of optimal control problems with minimal implementation effort.
@@
We can trace attempts to solve optimal control problems through probabilistic inference back to the early 2000s, beginning with path integral control [kappen2005path] and formulating stochastic optimal control as Kullback-Leibler (KL) divergence minimization [todorov2008general, kappen2012optimal]. It then extends to information-theoretic frameworks through Model Predictive Path Integral Control (MPPI) [williams2017information, williams2018information], and reaches connections to Reinforcement Learning (RL) through the general framework of Control as Inference [levine2018reinforcement]. Despite its theoretical elegance and growing impact in robotics and decision-making, its formulation is often scattered across different research communities, making it difficult for control theorists and robotic practitioners to obtain a coherent understanding. Moreover, existing tutorials tend to focus either on RL-based formulations [levine2018reinforcement] or on path integral optimal control [kazim2024recent]. While the former is commonly framed in terms of machine learning rather than MPC, the latter is primarily developed within control theory and does not always make its connection to variational or probabilistic inference explicit.
This paper aims to provide a comprehensive tutorial and survey on PI-MPC, serving as a guide for both theorists and practitioners. In the first half, we explain the theoretical foundations of PI-MPC, primarily based on [levine2018reinforcement, williams2018information, wang2021variational, abdolmaleki2018maximum]. Specifically, as illustrated in Fig. 1, we describe the approach in two steps. First, we derive the optimal control distribution from the optimal control problem and discuss its probabilistic interpretation and key characteristics. Second, we explain how to generate actions from the optimal control distribution using VI techniques, taking MPPI as a representative example [^1] We open-source a PyTorch-based MPPI implementation, available at https://github.com/kohonda/mppi_playground. In the latter half, we survey key challenges, representative methods, and recent advances in PI-MPC, including fundamental topics and cutting-edge research directions.
@@ 2.1 Definition of Optimal Control Problem
Here, $J$ is the cost function; $\mathbf{x}_{t}$ and $\mathbf{u}_{t}$ denote the state and control input at time $t$. We write the control sequence as $\mathbf{u}_{0:T-1}=[\mathbf{u}_{0},\dots,\mathbf{u}_{T-1}]$ and the corresponding state sequence as $\mathbf{x}_{0:T}=[\mathbf{x}_{0},\dots,\mathbf{x}_{T}]$. The dynamics are given by $\mathbf{f}$, and $\mathbf{x}_{0}$ is the observed initial state. Most optimal control problems also impose constraints on the state and the control input. In PI-MPC, input constraints can be incorporated into the dynamics, whereas state constraints are typically handled indirectly by embedding them in the cost function (e.g.,, via indicator functions). For details, refer to the experiments in [williams2018information] and Section 3.3.
Given $J$ and $\mathbf{f}$, the controller solves Eq. (1) online and applies the resulting action sequence $\mathbf{u}^{*}_{0:T-1}$. When $J$ and $\mathbf{f}$ admit simple structures (e.g., quadratic objectives and differentiable dynamics), numerical optimization can be efficient. However, when $J$ and/or $\mathbf{f}$ are highly nonlinear or do not provide reliable analytical derivatives, standard optimization methods can become difficult to apply in real time.
@@ 2.2 Motivations for Probabilistic Inference-Based MPC
Figure 2: Comparison of sampling-based MPC methods in a vehicle obstacle avoidance task [honda2024stein]. Random shooting (Left) selects a single best trajectory from samples, resulting in poor sample efficiency, whereas MPPI (Right) estimates an optimal control distribution and achieves significantly improved efficiency through distribution-level inference.
To overcome the sample efficiency challenges, we therefore introduce a probabilistic inference framework. Specifically, as illustrated in Fig. 1, we approximate the probabilistic distribution of optimal control inputs (hereafter referred to as the optimal control distribution) using samples and generate control input sequences from this distribution. Rather than searching for a single best sequence among random samples, we infer a distribution over good control sequences; this distribution-level view can substantially improve sample efficiency (Fig. 2). Beyond sample efficiency, casting control as distribution optimization offers additional benefits:
  • It naturally represents stochastic variability in behavior.
@@
  • The overall pipeline can be implemented end-to-end in a differentiable manner (e.g., in PyTorch).
@@ 2.3.1 Derivation of the Optimal Control Distribution
Here, $Z$ is the normalization constant, $\lambda$ is the temperature parameter, $J_{\tau}(\mathbf{u}_{0:T-1})$ is the trajectory cost obtained from the control input $\mathbf{u}_{0:T-1}$ via Eq. (1), and $p(\mathbf{u}_{0:T-1})$ represents the prior distribution over control inputs. Thus, the optimal control distribution is proportional to the product of a Boltzmann distribution and a prior distribution [^2] A similar expression appears in Direct Preference Optimization [rafailov2023direct], which has been used for aligning large language models.. The Boltzmann term assigns higher probability to lower-cost trajectories; we will discuss it in more detail later. We now derive this expression in three steps.
Graphical Model Representation of the Optimal Control Problem
@@
Approximating the Optimal Control Distribution via Kullback-Leibler Divergence Minimization
@@
are substituted into Eq. (4) [^3] Equation (5) can also be derived via ELBO minimization.. Equation (6) defines the Boltzmann distribution, which assigns higher probability to lower-cost trajectories. As illustrated in Fig. 4, the temperature parameter $\lambda$ controls its sharpness: smaller $\lambda$ concentrates mass near the minimum-cost region. Although the Boltzmann form can be motivated by maximum-entropy arguments, it is not the only choice; alternative likelihood models are possible [^4] For example, the Cross-Entropy Method uses a thresholded likelihood [botev2013cross].. We use the Boltzmann form because it leads to a simple and practical inference procedure.
@@
Computing the Optimal Control Distribution via the Method of Lagrange Multipliers
@@ 2.4.1 Influence of the Prior Distribution on the Optimal Control Distribution
As shown in Eq. (2), the optimal control distribution is expressed as the product of the Boltzmann distribution and the control input prior distribution. Thus, “optimal actions” are not simply those that minimize the cost in Eq. (1a); they also reflect the action prior (Fig. 5).
Reference [kappen2005path] illustrates this characteristic with the analogy of a “drunken spider crossing a pond.” Imagine a spider attempting to cross a pond via a narrow bridge. When sober (i.e., when the prior has small variance), the best strategy is to take the shortest route across the bridge. When drunk (i.e., when the prior variance is large), crossing the narrow bridge becomes risky, and a longer detour around the pond may be preferable. The lesson is that optimal actions depend not only on the cost function but also on the prior distribution.
Note that models that use Boltzmann distributions as optimal distributions for sample generation are called Energy-Based Models (EBM), well studied in machine learning [lecun2006tutorial]. Compared with standard EBMs, Eq. (2) includes an explicit action prior, which is induced by the system dynamics in Eq. (1b).
@@ 2.4.2 Decision-Making Through Symmetry Breaking
One intriguing phenomenon arising from the prior distribution is “symmetry breaking” [kappen2005path]. Symmetry breaking refers to the phenomenon where the number of modes in the optimal control distribution decreases over time as the noise level of the prior distribution diminishes.
@@ 2.4.3 Relationship Between Temperature Parameter and Optimal Control Distribution
The temperature parameter is a hyperparameter that must be tuned for each application. Smaller $\lambda$ encourages more aggressively cost-seeking behavior, which can improve performance but may reduce sample efficiency and robustness to disturbances. Larger $\lambda$ places more weight on the prior, which can improve sample efficiency and robustness but may become overly conservative. In many works, a fixed temperature parameter is manually tuned, though adaptive adjustment methods have also been proposed [watson2023inferring, pezzato2025sampling].
@@ 2.4.4 Factors Governing Sample Complexity
As we discuss in Section 2.5, in practice we approximate the optimal control distribution in Eq. (2) using Monte Carlo sampling. Accordingly, it is desirable for the target distribution to be well approximated with as few samples as possible. The number of samples needed to obtain a good approximation depends strongly on the distribution’s shape; this is often referred to as sample complexity.
@@ 2.5 Control Input Generation from the Optimal Control Distribution: The MPPI Example
The most widely used PI-MPC method is MPPI [williams2018information]. MPPI employs a multivariate Gaussian distribution with fixed covariance $\Sigma$ as the variational distribution. Specifically, setting $\theta=\mu_{0:T-1}=[\mu_{0},\dots,\mu_{T-1}]$ and
@@ 3 Survey of Probabilistic Inference-Based MPC Methods
As established in Section 2, the main design choices in PI-MPC reduce to two elements: the prior distribution $p(\mathbf{u}_{0:T-1})$, and the variational family $\pi_{\boldsymbol{\theta}}$. Each choice involves distinct trade-offs between solution quality, computational cost, and robustness. This section surveys the key challenges that arise when instantiating these choices for real-world robotic tasks, together with representative methods that address them.
@@ 3.1 Design of Prior Distributions
Figure 7: Effect of prior distribution mismatch and variance on the optimal control distribution. Larger deviation between the prior and the Boltzmann distribution deteriorates optimality, while increased prior variance mitigates this effect when mismatch is large. Priors located near the Boltzmann mode with small variance yield the most favorable performance, whereas distant low-variance priors produce the poorest outcome. The cost function is the same as in Fig. 4, and four Gaussian prior distributions $p(u)=\mathcal{N}(\mu,\sigma),\mu\in\{-3.0,0.0\}$ and $\sigma\in\{0.25,1.0\}$ are compared.
The prior distribution $p(\mathbf{u}_{0:T-1})$ serves a dual role in PI-MPC: it defines the sampling region from which candidate solutions are drawn, and it acts as a regularizer that shapes the notion of optimality through the KL penalty in Eq. (5). A poorly chosen prior wastes samples in low-probability regions of the optimal distribution, while an overly narrow prior risks missing the global optimum entirely. The principal lever for addressing this trade-off is to bring the prior close to the modes of the Boltzmann distribution while keeping its variance small enough for efficient sampling yet large enough to avoid premature convergence.
@@ 3.1.1 What is a Good Prior Distribution?
From an algorithmic standpoint, this picture is intuitive: in the MPPI algorithm (Section 2.5), the prior defines the search region from which all samples are drawn. Focusing this region around the modes of the Boltzmann distribution maximizes both sample efficiency and solution quality, which is supported by theoretical analyses [Yoon2022-jb, Tao2022-dj]. The methods reviewed below all aim, in different ways, to achieve this ideal: a prior with small variance positioned near an optimal mode (Fig. 7, bottom-left).
@@ 3.1.2 Balancing Exploration and Exploitation via Prior Distribution
One family of methods achieves this by replacing the Gaussian prior with a leptokurtic distribution, i.e., one with higher kurtosis and heavier tails. Mohamed et al.  [mohamed2022autonomous] used a normal–log-normal (NLN) mixture, and Vlahov et al.  [vlahov2024low] employed colored noise as the prior. Such distributions suppress high-frequency components while preserving low-frequency structure, yielding temporally smooth samples that concentrate around plausible trajectories without sacrificing coverage.
Temporal smoothness can also be imposed more directly. Post-processing filters [andrejev2025pi] and spline interpolation [bhardwaj2022storm] smooth samples after generation, while other methods redefine the action space in terms of action derivatives [kim2022smooth, lee2time], so that sampling in the new space inherently produces smooth trajectories. A more principled alternative is to embed temporal correlations into the inference problem itself: Gaussian Processes (GPs) naturally encode such correlations and can replace the temporally independent Gaussian prior used in standard MPPI [mukadam2018continuous, urain2022learning, watson2023inferring].

### v3

@@ Abstract
This paper presents a tutorial and survey on probabilistic inference-based model predictive control (PI-MPC) for robotics. PI-MPC reformulates finite-horizon optimal control as inference over an optimal control distribution expressed as a Boltzmann distribution weighted by a control prior, and generates actions through variational inference. In the tutorial part, we derive this formulation and explain action generation via variational inference, highlighting Model Predictive Path Integral (MPPI) control as a representative algorithm with a closed-form sampling update. In the survey part, we organize existing PI-MPC research around key design dimensions, including prior design, multi-modality, constraint handling, scalability, hardware acceleration, and theoretical analysis. This paper provides a unified conceptual perspective on PI-MPC and a practical entry point for robotics researchers and practitioners.
@@
journal: Nuclear Physics B
[…]
[label1]organization=Nagoya University, addressline=Furo-cho, Chikusa-ku, city=Nagoya, postcode=464-8603, state=Aichi, country=Japan
[…]
[label2]organization=CyberAgent AI Lab, addressline=Shibuya Scramble Square, city=Shibuya, postcode=150-6121, state=Tokyo, country=Japan
@@ 1 Introduction
Establishing a universal control theory applicable to diverse robots and tasks has long been a fundamental goal in robotics research. Model Predictive Control (MPC) offers an intuitive and general framework for robot control by optimizing actions while predicting future states over a finite horizon. To apply MPC to real-world robots and tasks, we may need to formulate complex optimal control problems that capture the intricacies of robot dynamics and task requirements. However, these formulations often exhibit strong nonlinearity, making them intractable for traditional numerical optimization methods widely familiar in the control field. In many robotic applications, what we need is not only strong theoretical guarantees, but also an easy-to-use optimization framework that can reliably solve a broad class of optimal control problems with minimal implementation effort.
@@
We can trace attempts to solve optimal control problems through probabilistic inference back to the early 2000s, beginning with path integral control (kappen2005path) and formulating stochastic optimal control as Kullback-Leibler (KL) divergence minimization (todorov2008general; kappen2012optimal). It then extends to information-theoretic frameworks through Model Predictive Path Integral Control (MPPI) (williams2017information; williams2018information), and reaches connections to Reinforcement Learning (RL) through the general framework of Control as Inference (levine2018reinforcement). Despite its theoretical elegance and growing impact in robotics and decision-making, its formulation is often scattered across different research communities, making it difficult for control theorists and robotic practitioners to obtain a coherent understanding. Moreover, existing tutorials tend to focus either on RL-based formulations (levine2018reinforcement) or on path integral optimal control (kazim2024recent). While the former is commonly framed in terms of machine learning rather than MPC, the latter is primarily developed within control theory and does not always make its connection to variational or probabilistic inference explicit.
This paper aims to provide a comprehensive tutorial and survey on PI-MPC, serving as a guide for both theorists and practitioners. In the first half, we explain the theoretical foundations of PI-MPC, primarily based on (levine2018reinforcement; williams2018information; wang2021variational; abdolmaleki2018maximum). Specifically, as illustrated in Fig. 1, we describe the approach in two steps. First, we derive the optimal control distribution from the optimal control problem and discuss its probabilistic interpretation and key characteristics. Second, we explain how to generate actions from the optimal control distribution using VI techniques, taking MPPI as a representative example [^1] We open-source a PyTorch-based MPPI implementation, available at https://github.com/kohonda/mppi_playground. In the latter half, we survey key challenges, representative methods, and recent advances in PI-MPC, including fundamental topics and cutting-edge research directions.
@@ 2.1 Definition of Optimal Control Problem
Here, $J$ is the cost function; $\mathbf{x}_{t}$ and $\mathbf{u}_{t}$ denote the state and control input at time $t$. We write the control sequence as $\mathbf{u}_{0:T-1}=[\mathbf{u}_{0},\dots,\mathbf{u}_{T-1}]$ and the corresponding state sequence as $\mathbf{x}_{0:T}=[\mathbf{x}_{0},\dots,\mathbf{x}_{T}]$. The dynamics are given by $\mathbf{f}$, and $\mathbf{x}_{0}$ is the observed initial state. Most optimal control problems also impose constraints on the state and the control input. In PI-MPC, input constraints can be incorporated into the dynamics, whereas state constraints are typically handled indirectly by embedding them in the cost function (e.g., via indicator functions). For details, refer to the experiments in (williams2018information) and Section 3.3.
Given $J$ and $\mathbf{f}$, the controller solves Eq. (1) online and applies the resulting action sequence $\mathbf{u}^{*}_{0:T-1}$. When $J$ and $\mathbf{f}$ admit simple structures (e.g., quadratic objectives and differentiable dynamics), numerical optimization can be efficient. However, when $J$ and/or $\mathbf{f}$ are highly nonlinear or non-differentiable, standard optimization methods can become difficult to apply in real time.
@@ 2.2 Motivations for Probabilistic Inference-Based MPC
(a) Random shooting
(b) MPPI
Figure 2: Comparison of sampling-based MPC methods in a vehicle obstacle avoidance task (honda2024stein). Random shooting selects a single best trajectory from samples, resulting in poor sample efficiency, whereas MPPI estimates an optimal control distribution and achieves significantly improved efficiency through distribution-level inference.
To overcome the sample efficiency challenges, we therefore introduce a probabilistic inference framework. Specifically, as illustrated in Fig. 1, we approximate the probabilistic distribution of optimal control inputs (hereafter referred to as the optimal control distribution) using samples to generate control input sequences from this distribution. Rather than searching for a single best sequence among random samples, we infer a distribution over good control sequences; this distribution-level view can substantially improve sample efficiency (Fig. 2). Beyond sample efficiency, casting control as distribution optimization offers additional benefits:
  • 1.
@@
  • 2.
    It naturally represents stochastic variability in behavior, which is beneficial for exploration and enables effective data augmentation, as described in Section 3.7.
  • 3.
    The overall pipeline can be implemented end-to-end in a differentiable manner (e.g., in PyTorch), enabling seamless integration with learning-based models and gradient-based optimization of both control and model parameters.
@@ 2.3.1 Derivation of the Optimal Control Distribution
Here, $Z$ is the normalization constant, $\lambda$ is the temperature parameter, $J_{\tau}(\mathbf{u}_{0:T-1})$ is the trajectory cost obtained from the control input $\mathbf{u}_{0:T-1}$ via Eq. (1), and $p(\mathbf{u}_{0:T-1})$ represents the prior distribution over control inputs. Thus, the optimal control distribution is proportional to the product of a Boltzmann distribution and a prior distribution [^2] A similar expression appears in Direct Preference Optimization (rafailov2023direct), which has been used for aligning large language models.. The Boltzmann term assigns higher probability to lower-cost trajectories; we will discuss it in more detail later. We now derive this expression in three steps.
Step 1: Graphical Model Representation of the Optimal Control Problem
@@
Step 2: Approximating the Optimal Control Distribution via Kullback-Leibler Divergence Minimization
@@
are substituted into Eq. (4) [^3] Equation (5) can also be derived via evidence lower bound minimization.. Equation (6) defines the Boltzmann distribution, which assigns higher probability to lower-cost trajectories. As illustrated in Fig. 4, the temperature parameter $\lambda$ controls its sharpness: a smaller $\lambda$ concentrates mass near the minimum-cost region. Although the Boltzmann form can be motivated by maximum-entropy arguments, it is not the only choice; alternative likelihood models are possible [^4] For example, the Cross-Entropy Method uses a thresholded likelihood (botev2013cross).. We use the Boltzmann form because it leads to a simple and practical inference procedure.
@@
Step 3: Computing the Optimal Control Distribution via the Method of Lagrange Multipliers
@@ 2.4.1 Influence of the Prior Distribution on the Optimal Control Distribution
As shown in Eq. (2), the optimal control distribution is expressed as the product of the Boltzmann distribution and the control input prior distribution. Thus, “optimal actions” are not simply those that minimize the cost in Eq. (1a); they also reflect the prior distribution (Fig. 5).
kappen2005path illustrates this characteristic with the analogy of a “drunken spider crossing a pond.” Imagine a spider attempting to cross a pond via a narrow bridge. When sober (i.e., when the prior has small variance), the best strategy is to take the shortest route across the bridge. When drunk (i.e., when the prior variance is large), crossing the narrow bridge becomes risky, and a longer detour around the pond may be preferable. The lesson is that optimal actions depend not only on the cost function but also on the prior distribution.
Note that models that use Boltzmann distributions as optimal distributions for sample generation are called Energy-Based Models (EBM), well studied in machine learning (lecun2006tutorial). Compared with standard EBMs, Eq. (2) includes an explicit action prior, which is induced by the system dynamics in Eq. (1b).
@@ 2.4.2 Decision-Making Through Symmetry Breaking
One intriguing phenomenon arising from the prior distribution is “symmetry breaking” (kappen2005path). Symmetry breaking refers to the phenomenon where the number of modes in the optimal control distribution decreases over time as the noise level of the prior distribution diminishes.
@@ 2.4.3 Relationship Between Temperature Parameter and Optimal Control Distribution
The temperature parameter is a hyperparameter that must be tuned for each application. Smaller $\lambda$ encourages more aggressive cost-seeking behavior, which can improve performance but may reduce sample efficiency and robustness to disturbances. Larger $\lambda$ places more weight on the prior, which can improve sample efficiency and robustness but may become overly conservative. In many works, a fixed temperature parameter is manually tuned, though adaptive adjustment methods have also been proposed (watson2023inferring; pezzato2025sampling).
@@ 2.4.4 Factors Governing Sample Complexity
As we discuss in Section 2.5, in practice, we approximate the optimal control distribution in Eq. (2) using Monte Carlo sampling. Accordingly, it is desirable for the target distribution to be well approximated with as few samples as possible. The number of samples needed to obtain a good approximation depends strongly on the distribution’s shape; this is often referred to as sample complexity.
@@ 2.5 Control Input Generation from the Optimal Control Distribution: The MPPI Example
The most widely used PI-MPC method is MPPI (williams2018information). MPPI employs a multivariate Gaussian distribution with fixed covariance $\Sigma$ as the variational distribution. Specifically, setting $\theta=\mu_{0:T-1}=[\mu_{0},\dots,\mu_{T-1}]$ and
@@ 3 Survey of Probabilistic Inference-Based MPC Methods
As established in Section 2.6, the main design choices in PI-MPC reduce to two elements: the prior distribution $p(\mathbf{u}_{0:T-1})$, and the variational family $\pi_{\boldsymbol{\theta}}$. Each choice involves distinct trade-offs between solution quality, computational cost, and robustness. This section surveys the key challenges that arise when instantiating these choices for real-world robotic tasks, together with representative methods that address them.
@@ 3.1 Design of Prior Distributions
The prior distribution $p(\mathbf{u}_{0:T-1})$ plays two conceptually equivalent roles in PI-MPC: it defines the sampling region from which candidate solutions are drawn, and it acts as a regularizer that shapes the notion of optimality through the KL penalty in Eq. (5). A poorly chosen prior wastes samples in low-probability regions of the optimal distribution, while an overly narrow prior risks missing the global optimum entirely. The principal lever for addressing this trade-off is to bring the prior close to the modes of the Boltzmann distribution while keeping its variance small enough for efficient sampling yet large enough to avoid premature convergence.
@@ 3.1.1 What is a Good Prior Distribution?
Figure 7: Effect of prior distribution mismatch and variance on the optimal control distribution. Larger deviation between the prior and the Boltzmann distribution deteriorates optimality, while increased prior variance mitigates this effect when the mismatch is large. Priors located near the Boltzmann mode with small variance yield the most favorable performance, whereas distant low-variance priors produce the poorest outcome. The cost function is the same as in Fig. 4, and four Gaussian prior distributions $p(u)=\mathcal{N}(\mu,\sigma),\mu\in\{-3.0,0.0\}$ and $\sigma\in\{0.25,1.0\}$ are compared.
@@
From an algorithmic standpoint, this picture is intuitive: in the MPPI algorithm (Section 2.5), the prior defines the search region from which all samples are drawn. Focusing this region around the modes of the Boltzmann distribution maximizes both sample efficiency and solution quality, which is supported by theoretical analyses (Yoon2022-jb; Tao2022-dj). The methods reviewed below all aim, in different ways, to achieve this ideal: a prior with small variance positioned near an optimal mode (Fig. 7, bottom-left).
@@ 3.1.2 Balancing Exploration and Exploitation via Prior Distribution
One family of methods achieves this by replacing the Gaussian prior with a leptokurtic distribution, i.e., one with higher kurtosis and heavier tails. mohamed2022autonomous used a normal–log-normal (NLN) mixture, and vlahov2024low employed colored noise as the prior. Such distributions suppress high-frequency components while preserving low-frequency structure, yielding temporally smooth samples that concentrate around plausible trajectories without sacrificing coverage.
Temporal smoothness can also be imposed more directly. Post-processing filters (andrejev2025pi) and spline interpolation (bhardwaj2022storm) smooth samples after generation, while other methods redefine the action space in terms of action derivatives (kim2022smooth; lee2time), so that sampling in the new space inherently produces smooth trajectories. A more principled alternative is to embed temporal correlations into the inference problem itself: Gaussian Processes (GPs) naturally encode such correlations and can replace the temporally independent Gaussian prior used in standard MPPI (mukadam2018continuous; urain2022learning; watson2023inferring).